# Curriculum Approximate Unlearning for Session-based Recommendation

### v1

@@ 5.5.2. Impact of unlearning data ratio
In this part, we investigate the performance of different models under varying amounts of unlearning data, as illustrated in the figure 3. GRU4Rec is chosen as the backbone recommender model and trained on Beauty dataset. We set the unlearning data ratio of the training data from 3% to 10%. We can observe that the variation in U-score for Retrain is minimal, as it represents the most ideal yet computationally expensive approach. CAU also exhibits a variation within 1% when the ratio changes, further indicating the greater stability and robustness of our method, particularly due to the contributions of the Pareto-Optimality solver and sampling strategy. In contrast, the stability of SISA and SRU decreases as the amount of unlearning data increases, with SISA dropping by 8% and SRU by 4%.
@@ 6. Conclusion
In this paper, we propose CAU, a curriculum-based approximate unlearning framework specifically designed for session-based recommendation systems. By formulating unlearning as a multi-objective optimization problem and seeking Pareto-Optimal solutions, CAU effectively balances the trade-off between forgetting targeted samples and preserving overall recommendation performance. Furthermore, the curriculum learning strategy, guided by the proposed gradient and embedding unlearning difficulty scores, along with the hard-sampling and soft-sampling methods, enables a more efficient and effective unlearning process. Extensive experiments on multiple benchmark datasets demonstrate the superiority of CAU over existing methods, confirming its practical value and robustness in handling multiple unlearning requests.
Ethical considerations
Our proposed method for curriculum approximate unlearning in session-based recommendation systems aims to remove the influence of specific user data efficiently. While this can enhance user privacy and control over personal data, potential negative societal impacts exist. For example, improper unlearning could inadvertently degrade recommendation quality for certain users. However, these risks are minimal and can be effectively managed.

### v2

@@ 5.5.2. Impact of unlearning data ratio
Finally, we investigate the performance of different models under varying amounts of unlearning data, as illustrated in Figure 3. GRU4Rec is chosen as the backbone recommender model and trained on the Beauty dataset. We set the unlearning data ratio of the training data from 3% to 10%. We can observe that the variation in U-score for Retrain is minimal, as it represents the most ideal yet computationally expensive approach. CAU also exhibits a variation within 1% when the ratio changes, further indicating the greater stability and robustness of our method, particularly due to the contributions of the Pareto-optimality solver and sampling strategy. In contrast, the stability of SISA and SRU decreases as the amount of unlearning data increases, with SISA dropping by 8% and SRU by 4%.
@@ 6. Conclusion
In this paper, we have proposed CAU, a curriculum-based approximate unlearning framework specifically designed for session-based recommendation systems. By formulating unlearning as a multi-objective optimization problem and seeking Pareto-Optimal solutions, CAU effectively balances the trade-off between forgetting targeted samples and preserving overall recommendation performance. Furthermore, the curriculum learning strategy, guided by the proposed gradient and embedding unlearning difficulty scores, along with the hard-sampling and soft-sampling methods, enables a more efficient and effective unlearning process. Extensive experiments on multiple benchmark datasets demonstrate the superiority of CAU over existing methods, confirming its practical value and robustness in handling multiple unlearning requests.
Nevertheless, certain limitations still persist in the current study. While our curriculum strategy effectively prioritizes unlearning batches, the long-term impact on model fairness across heterogeneous session lengths remains understudied. Regarding potential directions for future work, we plan to further validate CAU’s generalizability by adapting it to diverse recommendation architectures beyond session-based recommendation, such as graph-based or cross-domain recommendation models.